# RING-NeRF: A Versatile Architecture based on Residual Implicit Neural Grids

### v1

@@ 3.3 Distance-Aware forward mapping
Cone Casting. In order to accommodate the scene rendering process to the variation of the observation distance, we introduce a distance-aware forward mapping mechanism. Similarly to the cone casting of [6], it relies on assigning to each sample a latent feature whose LOD is inversely proportional to the sample-camera distance. Moreover, the continuous LOD property of our architecture (see eq. 2) assures smooth feature transitions as the distance to the camera grows.
To compute the LOD $L\in\mathbb{R}^{+}$ corresponding to a sample located at distance $d$, let’s first define $c$ as the length of a pixel’s side at distance $1$ of the camera, $b$ as the resolution of the coarsest grid $G_{0}$ of the hierarchy, and $f$ the growth factor between the resolution of two consecutive grids in the hierarchy such that the resolution of the grid $n$ is $f^{n}b$. Based on these conventions, the size of a cell side of the grid corresponding to the LOD $L$ is equal to $1/f^{L}b$.
Then, we define the LOD $L$ corresponding to the distance d as the LOD of the virtual grid [^3] A virtual grid of LOD L corresponds to a grid of resolution $f^{L}b$ that is not explicitly stored in memory but whose elements can be computed from other grids . whose cell volume $(d.c)^{3}$ is equal to the volume of our projected pixel at distance $d$. Hence :
$(d.c)^{3}=(\frac{1}{f^{L}b})^{3}\iff L=\frac{-log(d.c.b)}{log(f)}$ (3)
@@
Forward Mapping. As illustrated in Figure 2, we use the determined LOD to compute a distance-dependent weighted sum of the existing features and feed them to our decoder, which transforms them into density and radiance.
3.4 Continuous Coarse-To-Fine Optimization
@@
Recent works proposed to use coarse-to-fine optimization to improve the stability of NeRF models, especially when facing harder setups, including with few images [19] and surface-based models [10]. It consists of optimizing progressively the different LOD of the model, from coarse levels to the most precise ones. The goal of this progressive optimization is to avoid the shape-radiance ambiguity [22] by introducing a strong regularization through LOD restriction, then relaxing progressively this regularization once the coarse geometry of the scene is reconstructed to recover the details of the scene.
@@
Since our scene representation relies on a multi-LOD mapping function between the scene space and the latent space, the coarse-to-fine reconstruction process consists of estimating progressively the LODs of the mapping function from the coarsest to the finest ones. In practice, it implies, during the cone casting, to clamp the samples’ LOD up to a maximal LOD $l$, the grids of level $l$ and above being set to zero and not optimized. Moreover, since our architecture provides continuous LOD, this optimization can be achieved continuously in the LOD space by using a linear scheduler $l=l_{0}+\frac{n}{n_{ctf}}$ with $n$ the current epoch, $n_{ctf}$ a hyperparameter describing the speed of the process and $l_{0}$ defining the number of used grids at initialization; up to a specified maximum resolution.
@@
While this strategy, consisting of filling the grids with zeros and optimizing progressively an increasing number of levels of the hierarchical grid, was already used with I-NGP-like architectures [10], the underlying process is different. Indeed, for solutions based on the concatenation of features, keeping grids values to zero implies that some dimensions of the decoder’s latent space are not supervised. When a grid starts to be optimized, those unsupervised dimensions are suddenly used. The weights of the decoder thus need to be refined, with a global effect on the whole scene. On the opposite, our solution keeps supervising all the dimensions of the decoder latent space, and when a grid gets optimized, it only implies more degrees of freedom to define the mapping function between the scene space and the decoder latent space. It means that the scene reconstruction can be refined even without modifying the decoder weight, as we demonstrate experimentally in section 4.5.
Moreover, we also decide to progressively and continuously unmask higher frequencies of the color-inducing view direction encoding. In order to do this, we use Spherical Harmonics [5] and progressively release the encoding using the linear scheduler previously described.
@@ 4.1 Implementation
Our model is based on the PyTorch framework Nerfstudio [18]. We build upon its core method named Nerfacto, which combines ideas from several papers for fast and qualitative renders of unbounded complex scenes. It is in fact an improved version of grid-based Instant-NGP [13] with various elements of MipNeRF-360 [2]: mostly hierarchical HashGrids, scene contraction and proposal samplers. This makes it an accessible baseline with a state-of-the-art quality/time ratio.
Configuration details, further experiments analysis, and illustrations can be found in the supplementary materials.
@@ 4.2 Novel View Synthesis and Anti-Aliasing
This experiment aims to evaluate the reconstruction quality through the ability to synthesize viewpoints that were not supervised during the reconstruction. These new viewpoints will differ from the angle of observation, but also the distance of observation. The latter is particularly important since a reconstruction or a rendering process that does not take correctly into account the distance of observation leads to rendering artifacts, from over-contrasted rendering to aliasing phenomenon. The challenge is then to avoid these artifacts while keeping the reconstruction and rendering process as fast as possible.
@@
Dataset. The evaluation relies on the dataset introduced by Mip-NeRF-360 [2]. This dataset is composed of 9 scenes, each containing both a central area and complex background in both inside and outside setups. Since the trajectory keeps a constant distance with respect to the central part, each viewpoint is represented with a pyramid of 4 different image resolutions to simulate a variation of the distance of observation combined with a change of the camera FOV, following the Mip-NeRF [1] original anti-aliasing evaluation pipeline.
Algorithms. We compare our solution against several usual NeRF baselines, both coordinate-based and grid-based, using NerfStudio implementation for the two main baselines Nerfacto and a reproduction of ZipNeRF [3], the source code of the latter being not publicly available. For NeRF, Mip-NeRF and Mip-NeRF-360, the reported performances are extracted from the ZipNeRF publication. Purely object-centric solutions, such as Tri MipRF [6], were not considered since out of the scope.

### v3

@@ 3.3 Distance-Aware Forward Mapping
Cone Casting in contracted space. In order to accommodate the scene rendering process to the variation of observation distances, we introduce a distance-aware forward mapping mechanism. Similarly to the cone casting of [7], it relies on assigning to each sample a latent feature whose LOD is inversely proportional to the sample-camera distance. However, unlike [7, 21], our model relies on the use of space contraction to allow the reconstruction of unbounded scenes.
To define the LOD of a sample at distance $d$ of the camera, we first compute the associated volume of the cone cast pixel cube in the world space coordinates (see fig. 2). Assuming the pixel is a square of size $c$ at distance $1$ of the camera ($c$ depends on the image resolution and camera’s FOV), the volume is thus $V=(dc)^{3}$. To take into account the space contraction, we then proceed to contract the volume. Denoting $J$ the Jacobian of the contraction function at the sample’s location $p$, $V_{contract}=Vdet(J(p))$. In practice, we compute the analytical derivation of the contraction function depending on $p$. More details can be found in the supplementary materials.
Assuming the $N$-th feature grid’s resolution in our hierarchy can be written as $f^{N}b$ with $b$ the resolution of the grid $\delta_{0}$ and $f$ the growth factor, finding the appropriate LOD $L\in\mathbb{R}^{+}$ means finding a virtual grid [^2] A virtual grid of LOD $L\in\mathbb{R}^{+}$ corresponds to a grid of resolution $f^{L}b$ that is not explicitly stored in memory but whose elements can be computed from other grids. of resolution $f^{L}b$ whose cell’s volume is equal to the previously computed volume (in contracted space). Because we are working in the contracted space of size $1$, the volume of a cell of the virtual grid of LOD $L$ is $(\frac{1}{f^{L}b})^{3}$. The LOD $L$ associated to one sample in the contracted space is thus given by:
$(dc)^{3}\operatorname{det}(J(p))=\left(\frac{1}{f^{L}b}\right)^{3}\iff L=-\frac{\log\left(dcb\sqrt[3]{\operatorname{det}(J(p))}\right)}{\log(f)}$ (3)
Note that this process is close to ZipNeRF [4] and VR-NeRF [24]. However, the first one rather derives a contracted scale factor from its Gaussian samples while the latter directly computes the LOD in the contracted space, which can be considered an approximation of our computation.
@@
Forward mapping. As illustrated in Figure 2, we use the determined LOD to compute a distance-dependent weighted sum of the features, which is fed to our decoder and transformed into density (or SDF) and radiance.
3.4 Continuous Coarse-To-Fine and Resolution Extensibility
@@
Recent works proposed to use coarse-to-fine optimization to improve the stability of NeRF models, especially when facing more challenging setups, including with few images [25] and surface-based models [11]. It consists of optimizing progressively the different LOD of the representation, from coarse levels to the most precise ones. The goal of this progressive optimization is to avoid the shape-radiance ambiguity [28] by introducing a strong regularization through LOD restriction, then relaxing progressively this regularization once the coarse geometry of the scene is reconstructed to recover the details of the scene.
@@
The coarse-to-fine reconstruction process of RING-NeRF consists of estimating progressively the LOD of the mapping function from the coarsest to the finest ones. In practice, it implies, during the cone casting, to clamp the samples’ LOD up to a maximal LOD $l$, the grids of level $l$ and above being set to zero and not optimized. Moreover, since our architecture provides continuous LOD, the coarse-to-fine optimization can be achieved continuously in the LOD space by using a linear scheduler $l=(l_{0}+\frac{n}{n_{ctf}})\in\mathbb{R}^{+}$ with $n$ the current epoch, $n_{ctf}$ a hyperparameter describing the speed of the process and $l_{0}$ defining the number of used grids at initialization; up to a specified maximum resolution.
@@
Furthermore, the RING-NeRF architecture is more adapted than I-NGP-based architectures [11] for coarse-to-fine training. Indeed, for solutions based on the concatenation of features, keeping grid values to zero implies that some dimensions of the decoder’s latent space are not supervised. When a grid starts to be optimized, those unsupervised dimensions are suddenly used. The weights of the decoder thus need to be refined, with a global effect on the whole scene. On the opposite, our solution keeps supervising all the dimensions of the decoder latent space, and when a new grid gets optimized, it only implies more degrees of freedom to define the mapping function between the scene space and the decoder latent space. This also means that our scene reconstruction can be refined by adding dynamically new grid levels without modifying the decoder’s weights or previously trained grids, as we demonstrate experimentally in section 4.5. This resolution extensibility property opens the path to adaptive resolution models, where the precision used to describe an area depends on the details needed, to optimize efficiency both in memory consumption and training duration.
@@ 4 Experiments
In these experiments, we intend to highlight the versatility of RING-NeRF by evaluating it on several tasks. After introducing implementation details in section 4.1, we evaluate our model on novel view synthesis with changes in observation distances (sec. 4.2). Then, we explore how robust is RING-NeRF first with few view reconstruction (sec. 4.3) and then without scene-specific initialization for SDF reconstruction (sec. 4.4). Finally, we demonstrate the capacity of our architecture to perform LOD extensibility, as a first step towards adaptive reconstruction (sec. 4.5).
@@ 4.1 Implementation
Our model is based on the PyTorch framework Nerfstudio [20]. We build upon its core method named Nerfacto, which combines ideas from several papers for fast and qualitative renders of unbounded complex scenes. This makes it an accessible baseline with a state-of-the-art quality/time ratio. Because NeRF pipelines contain a high number of small but decisive choices of implementation (eg. some frameworks choose to train their models image by image while Nerfstudio jointly and randomly samples across all images), we decided to use as much as possible Nerfstudio-based baselines for fairer comparisons. All of these models are trained on one Nvidia-A100 GPU. The reported times correspond to the approximated training duration of the models. Configuration details, further experiments and ablatives on our contributions can be found in the supplementary materials.
Examples of image renderings at 1/8th resolution from models solely trained with the full resolution images. RING-NeRF is the only method capable of producing coherent aliasing-free renderings thanks to its LOD inductive bias.
@@ 4.2 Novel View Synthesis and Anti-Aliasing
This experiment aims to evaluate the reconstruction quality through the ability to synthesize viewpoints that are not supervised during the reconstruction. These new viewpoints differ from the angle of observation, but also from the distance of observation. The latter is particularly important since a reconstruction or a rendering process that does not take correctly into account the distance of observation leads to artefacts, from over-contrasted rendering to aliasing phenomenon. The challenge is to avoid these artefacts while keeping the reconstruction and rendering process as fast as possible.
@@
Dataset. The evaluation relies on the dataset introduced by Mip-NeRF-360 [3]. This dataset is composed of 9 scenes, each containing both a central area and complex background in both inside and outside setups. Since the trajectory keeps a constant distance to the central part, each viewpoint is represented with a pyramid of 4 different image resolutions to simulate a variation of the distance of observation combined with a change of the camera FOV, following the Mip-NeRF [2] original anti-aliasing evaluation pipeline.
Algorithms. We compare our solution against several grid-based NeRF baselines, both with (PyNeRF [21] and Zip-NeRF [4]) and without anti-aliasing processing (Nerfacto), using their Nerfstudio implementations.